93 3 pts Question2 Travel Agency A travel agency priced tickets between Chicago and Orlando. The passenger was to fly coach class round-trip, staying seven days. Calls were made directly to airlines and to travel agents with the following results. Note that the data reflects round-trip airfare. $227.00 $345.00 $600.00 $228.00 $428.00 $603.00 345.00 $229.00 $605.00 $339.00 $279.00 $407.00 Find the standard deviation of the quoted fare. (Round to the nearest 2 decimal places if needed). D Question 3 3 pts Shopping Duration Distribution Recently, a store manager tracked the time customers spent in the store from the time they took a number until they left. A sample of 12 customers was selected and the following data was recorded; where the frequency is the number of customers and 26 MacBook Ai
Answer #1

2. To find standard deviation we need to first find mean=\mu=\frac{\sum x}{n}=386.25

Create the following table.

data data-mean (data - mean)2
227 -159.25 25360.5625
345 -41.25 1701.5625
600 213.75 45689.0625
228 -158.25 25043.0625
428 41.75 1743.0625
603 216.75 46980.5625
345 -41.25 1701.5625
229 -157.25 24727.5625
605 218.75 47851.5625
339 -47.25 2232.5625
279 -107.25 11502.5625
407 20.75 430.5625

Find the sum of numbers in the last column to get. \sum (xi-\overline{x})^2=234964.25

Hence \sigma=\sqrt{\frac{\sum (xi-\overline{x})^2}{n-1}}=\sqrt{\frac{234964.25}{12-1}}=146.1519